Write down the equation of the line AB through (3,2) and perpendicular to the line 2y = 3x + 5.

Answer by mananth(16946)

2 y = 3 x + 5
Divide by 2
y = 1 1/ 2 x + 2 1/2
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1 1/2

The slope of a line perpendicular to the above line will be the negative reciprocal - 2/3
Because m1*m2 =-1
The slope of the required line will be - 2/3

m= - 2/3 ,point ( 3 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = -2 + b
b= 4
m= - 2/3
The required equation is y = - 2/ 3 x + 4